# Self-Admitted Technical Debt in LLM Software: An Empirical Comparison with ML and Non-ML Software

### v2

@@ Abstract
Self-admitted technical debt (SATD), referring to comments flagged by developers that explicitly acknowledge suboptimal code or incomplete functionality, has received extensive attention in machine learning (ML) and traditional (Non-ML) software. However, little is known about how SATD manifests and evolves in contemporary Large Language Model (LLM)-based systems, whose architectures, workflows, and dependencies differ fundamentally from both traditional and pre-LLM ML software. In this paper, we conduct the first empirical study of SATD in the LLM era, replicating and extending prior work on ML technical debt to modern LLM-based systems. We compare SATD prevalence across LLM, ML, and non-ML repositories across a total of 477 repositories (159 per category). We perform survival analysis of SATD introduction and removal to understand the dynamics of technical debt across different development paradigms. Surprisingly, despite their architectural complexity, our results reveal that LLM repositories accumulate SATD at similar rates to ML systems (3.95% vs. 4.10%). However, we observe that LLM repositories remain debt-free 2.4x longer than ML repositories (a median of 492 days vs. 204 days), and then start to accumulate technical debt rapidly. Moreover, our qualitative analysis of 377 SATD instances reveals three new forms of technical debt unique to LLM-based development that have not been reported in prior research: Model-Stack Workaround Debt, Model Dependency Debt, and Performance Optimization Debt. Finally, by mapping SATD to stages of the LLM development pipeline, we observe that debt concentrates significantly higher in the pretraining and deployment stages.
@@ I Introduction
Technical debt, the hidden cost of shortcuts in software development, has challenged software development since Cunningham first coined the metaphor in 1992 [6]. When developers write comments such as “TODO: fix this properly” or “HACK: temporary workaround”, they explicitly acknowledge these shortcuts. Such self-admitted technical debt (SATD) comments serve as signals, marking where future maintenance is likely needed [29, 22]. Prior work has shown that SATD predicts maintenance challenges [20], guides refactoring decisions [38], and reveals developer concerns about code quality [37].
@@
As machine learning (ML) became widespread, researchers discovered new SATD patterns, including experimental pipelines and data dependencies, unlike traditional (non-ML) software, where technical debt primarily stems from code structure and design violations [31, 4]. Since ChatGPT’s release in November 2022, Large Language Model (LLM) development has introduced fundamentally new architectural patterns that differ from both traditional and classical ML systems [28]. Modern LLM applications embed logic in prompts, use vector stores for Retrieval-Augmented Generation (RAG), coordinate multi-model workflows via agent frameworks, and depend on volatile external APIs [19, 8, 9]. These architectural shifts have introduced new forms of technical debt. For example, at the prompt layer, minor wording changes can break entire workflows [25], while orchestration frameworks may obscure critical business logic across distributed components [13]. At the infrastructure layer, token-based pricing can introduce economic debt that might compound with scale [30], and dependency on rapidly changing external APIs can create model lock-in risks [12]. Though practitioners encounter these issues daily [39], systematic empirical evidence remains limited.
@@
Most existing SATD research is based on data collected before the LLM revolution, with datasets [29, 37, 4, 27] that do not capture modern LLM architectural patterns. For example, Bhatia et al. conducted an empirical analysis of SATD in ML and non ML systems and provided a taxonomy of such debts without considering LLM-based systems [4]. Also, even recent work on LLM technical debt [1] examined only single-model prompt invocation and does not capture modern multi-agent, multi-model, or RAG-based workflows, nor does it examine SATD evolution over time or compare LLM against ML and non-ML software systems. In this paper, we conduct the first empirical study of SATD in the LLM era. We replicate and extend Bhatia et al.’s study [4] by constructing a new curated dataset of 159 LLM repositories (developed since November 2022) and comparing them against 159 ML and 159 non-ML repositories from Bhatia et al.’s dataset, using the latest updates to their metadata, commit history, and code comments. We analyze over five million comment events to identify patterns in the introduction, persistence, and removal of SATD across various development paradigms.
@@
Our experiments show that, despite their architectural complexity, LLM repositories accumulate SATD at similar rates to ML ones (3.95% vs. 4.10%) but remain debt-free $2.4x$ longer than ML repositories (a median of 492 days vs. 204 days) before debt starts to accumulate rapidly. We identify three new SATD categories specific to LLMs, accounting for 18% of the classified debt: Model-Stack Workaround Debt (7.4%), Model Dependency Debt (3.7%), and Performance Optimization Debt (4.0%). When mapped to LLM development pipeline [14], we observe that debt is most concentrated in Deployment and Monitoring (30%) and Pretraining (23%) stages. Our analysis also shows that SATD is rarely removed once introduced (removal rates below 5%), with developers tending to clean more recent, localized debt while leaving foundational issues intact. The first SATD introduced into any file proves nearly permanent (only 4.2% ever removed in LLM repositories) despite 49.1% overall comment-level removal. Our predictive modeling achieves 82% accuracy for LLM debt versus 96% for ML debt. This gap suggests that while ML debt remains highly predictable from traditional software metrics (confirming Bhatia et al.), LLM debt involves new factors (prompt engineering, RAG orchestration, API volatility) that conventional commit-level features cannot fully capture.
This work makes the following contributions:
@@
    We present the first empirical comparison of SATD across LLM, ML, and traditional software, showing that LLM development has unique debt accumulation patterns.
@@
    We extend the existing SATD taxonomy [4] by identifying three new technical debt types specific to LLM systems.
@@
    We provide insights on where debt accumulates (deployment >> pretraining >> infrastructure), predictors of persistence, and how debt removal differs across paradigms.
@@
    We release our complete replication package [2], including analysis scripts, extended SATD detector, and a curated dataset of 159 LLM repositories with full git histories for replication and future studies.
@@ II Related Work
Technical Debt in Machine Learning Systems. ML systems introduce forms of technical debt that go beyond those found in traditional systems. Sculley et al. first highlighted how ML systems accumulate data dependency debt, configuration debt, and boundary erosion, noting that seemingly small design shortcuts can propagate unpredictably through data pipelines and model behavior [31]. Building on this theoretical framework, Bhatia et al. [4] conducted a landmark empirical study comparing SATD in 318 ML and 318 non-ML repositories, finding that ML systems contain twice the SATD density of traditional software and exhibit distinct temporal patterns, introducing debt $140x$ faster but also removing it $3.7x$ faster. Their work established the empirical foundation for understanding ML technical debt at scale. O’Brien et al. [27] further expanded this understanding by identifying 23 distinct types of ML-specific technical debt through qualitative analysis. Together, these studies underscore that ML systems carry structural, data-centric, and operational debt patterns that differ significantly from traditional systems, with Bhatia et al.’s quantitative findings providing the empirical basis that motivates our replication and extension to LLM systems.
Emerging Technical Debt in LLM-Based Development. Empirical research on LLM-specific SATD is still nascent. Aljohani and Do [1] examined SATD in early LLM API usage patterns and found that prompt design rather than code structure is a primary source of LLM-specific TD. However, their focus on single-model prompt invocation does not capture modern multi-agent, multi-model, or RAG-based workflows, nor does it examine SATD evolution over time.
@@ III Research Questions
This study replicates and extends Bhatia et al.’s [4] analysis of SATD in ML systems to explore how SATD manifests in LLM systems. We examine whether architectural shifts in LLM development, such as prompt engineering, external APIs, and multi-model orchestration, affect technical debt patterns compared to traditional ML and non-ML software. We therefore study and compare the prevalence of SATD in LLM, ML and non-ML systems. We also study their types, stages and survival in such systems. We structure our study around five research questions that mirror and extend those of Bhatia et al. [4], as follows.
@@
RQ2. What are the different types of SATD in LLM-based systems? The original study extended Bavota and Russo’s SATD taxonomy to include ML-specific categories like configuration debt and inadequate testing. LLM development introduces new components absent in classical ML, such as RAG systems, vector databases, human-in-the-Loop workflows, prompt templates, embedding models, token-cost optimization, context-window management, and agent orchestration. These elements may create new debt categories within complex multi-model pipelines. We explore whether LLM repositories exhibit distinct SATD patterns that require further taxonomy extensions beyond pre-LLM ML systems.
RQ3. Which stages of the LLM pipeline are more prone to SATD? Bhatia et al. [4] found that model building and data preprocessing stages accumulate the most debt in classical ML pipelines. We investigate whether this holds for LLM systems or if debt instead concentrates in new architectural components, such as prompt orchestration, agentic logic, and RAG/vector database integration, or in LLM-specific stages like pretraining, fine-tuning, or deployment and monitoring.
@@ IV-A Replication Scope and Goals
To study SATD in LLM systems, we replicated and extended Bhatia et al.’s study of SATD in ML systems [4]. Their original analysis examined SATD in 318 ML and 318 non-ML Python projects using SATD detection and survival analysis to characterize SATD prevalence, types, and lifecycle. Our goals are twofold: (1) Replication: Reproduce their main quantitative findings on SATD prevalence and survival in ML and non-ML repositories using their replication package, detector, and analysis pipeline, adapted to current software versions, as the original replication package required updates to deprecated library dependencies. (2) Extension: Extend the study to contemporary, post-ChatGPT LLM repositories by constructing a curated dataset of LLM projects and applying the same SATD detection and survival analysis, as well as a refined SATD taxonomy tailored to LLM-specific development practices.
Wherever possible, we adhere to the original study’s design decisions, including project curation criteria, comment extraction, SATD detection, and statistical tests, while clearly documenting deviations, such as temporal window, LLM-specific search queries, and taxonomy extensions.
@@ IV-B LLM Dataset Curation
Similar to Bhatia et al. [4], we collected balanced repository sets with complete project evolution history via GitHub API to retrieve active systems, following established MSR practices [15, 26]. We constructed 41 search queries combining LLM-specific terms (“chatbot”, “gpt”, “openai”, “anthropic”), frameworks (“langchain”, “llama-index”), and architectural markers (“rag”, “embedding”, “agent”) with quality filters adapted from Bhatia et al.: (C1) non-forked repositories, (C2) $\geq 5\,$ Python files, (C3)>1>1 month development history, and (C4) LLM-specific libraries present. Data collection was completed in early November 2025. A multi-signal classification pipeline confirmed LLM relevance through (1) README semantics, (2) metadata/topic labels, and (3) dependency inspection. We organized 159 LLM repositories into a two-layer taxonomy: Infrastructure/Tools (n=101) and End-User Applications (n=58), with functional tags including Evaluation/Testing, Agentic, Training/Fine-tuning, Serving/Inference, RAG, General Chatbot, and Prompt Engineering. Our dataset targets post-November 2022 repositories to capture contemporary LLM constructs (e.g., prompt engineering, RAG, agentic orchestration, API-driven invocation), distinct from Bhatia et al.’s pre-LLM dataset (2015-2021), enabling faithful SATD replication within the LLM era.
@@ IV-C ML and Non-ML Repository Selection
Though Bhatia et al. [4] analyzed 318 ML and 318 non-ML Python projects, we sampled 159 repositories per category to match our LLM dataset as in Bavota and Russo [3]. Their work, which forms the basis for SATD classification in both Bhatia et al.’s and our study, showed that 159 repositories provide sufficient statistical power for SATD analysis while remaining computationally feasible for manual classification. This sample size ensures balanced comparison across LLM, ML, and non-ML repositories, without overrepresenting any type. For ML repositories, we randomly selected 159 projects from Bhatia et al.’s replication package [^1] https://drive.google.com/drive/folders/1n-gwAxFANS-PPewnk1Qgmnwd0zMAr7Yb, ensuring coverage across their six domains (Image Processing, Natural Language Processing (NLP), Audio Processing, Autonomous Gameplay, Self-Driving Cars, Other ML). For non-ML repositories, we replicated their GitHub API methodology using the keywords ”python projects” and ”server” with identical filtering criteria (non-forked, $\geq$5 Python files, >>1 month history, no ML components), yielding 159 repositories that matched the ML dataset characteristics.
@@ IV-D SATD Detection Pipeline
We employed Liu et al.’s NLP-based SATD detector [21], a state-of-the-art SATD classification tool that achieved 0.82 F1-score in prior evaluations. The detection process involved three steps: (1) extracting Python comments with Comment-Parser, (2) applying the pre-trained classifier to identify SATD instances, and (3) tracking temporal metadata for survival analysis. For RQ1, we analyze SATD prevalence in the current codebase (502,592 unique comments in LLM repositories). For RQ4-RQ5, we track 3,713,429 comment events (additions and removals) across all commits in LLM repositories, enabling both cross-sectional prevalence assessment and longitudinal lifecycle tracking.
@@ IV-E SATD Sampling and Classification Procedure
To address RQ2, we selected a representative subset of SATD comments from the dataset obtained in RQ1. Using stratified sampling across five domains with a 95% confidence level and 5% margin of error, we obtained a random sample of 377 SATD instances. This ensures balanced domain representation and reduces bias toward any single application area. Before coding the sampled comments, two co-authors jointly examined an initial set of 70 SATD instances (distinct from the 377 used in the final analysis) to align on the classification scheme. Coding followed a consensus-based approach and a card-sorting process [34]: both authors discussed each instance, resolving uncertainties through collaborative negotiation. We report raw agreement: 60 of the 70 instances (85.7%) were coded without disagreement, 10 instances (14.3%) required further discussion before reaching consensus. Through this iterative discussion, the two co-authors also identified patterns that did not fit existing categories and collectively defined three new LLM-specific debt types. This collaborative approach ensured consistent interpretation while allowing for emergent category discovery, which is particularly important given the novel characteristics of LLM-related technical debt.
After that, we categorized the 377 sampled SATD comments using a card-sorting approach guided by the extended taxonomy proposed by Bavota and Russo [3] [^2] Bhatia et al. [4] further adapted this taxonomy by adding Configuration Debt and Inadequate Testing, both of which we incorporate in our analysis.. Card sorting facilitated iterative comparison, refinement of category boundaries, and consistent labeling based on the initial joint examination. During coding, we encountered SATD comments that contained only the keyword “TODO” with no or little explanatory description of the intended action. We labeled such instances as Undefined, while those lacking context for technical debt were labeled as Unknown. In total, 44% of the sampled SATD comments were either false positives or belonged to the Undefined or Unknown categories. Excluding these, we classified 143 relevant SATD instances for RQ2. If an SATD instance did not fit the existing taxonomy, we introduced a new category, similar to Bhatia et al.’s extension of Hindle et al.’s taxonomy for ML systems [11].

### v3

@@ Abstract
Self-admitted technical debt (SATD), referring to comments in which developers explicitly acknowledge suboptimal code or incomplete functionality, has received extensive attention in machine learning (ML) and traditional (Non-ML) software. However, little is known about how SATD manifests and evolves in contemporary Large Language Model (LLM)-based systems, whose architectures, workflows, and dependencies differ fundamentally from both traditional and pre-LLM ML software. In this paper, we conduct the first empirical study of SATD in the LLM era, replicating and extending prior work on ML technical debt to modern LLM-based systems. We compare SATD prevalence across LLM, ML, and non-ML repositories across a total of 477 repositories (159 per category). We perform survival analysis of SATD introduction and removal to understand the dynamics of technical debt across different development paradigms. Surprisingly, despite their architectural complexity, our results reveal that LLM repositories accumulate SATD at similar rates to ML systems (3.95% vs. 4.10%). However, we observe that LLM repositories remain debt-free 2.4x longer than ML repositories (a median of 492 days vs. 204 days), and then start to accumulate technical debt rapidly. Moreover, our qualitative analysis of 377 SATD instances reveals three new forms of technical debt unique to LLM-based development that have not been reported in prior research: Model-Stack Workaround Debt, Model Dependency Debt, and Performance Optimization Debt. Finally, by mapping SATD to the stages of LLM development pipeline, we observe that debt concentrates significantly higher in the deployment and pretraining stages.
@@ I Introduction
Technical debt, the hidden cost of shortcuts in software development, has challenged software development since Cunningham first coined the metaphor in 1992 [5]. When developers write comments such as “TODO: fix this properly” or “HACK: temporary workaround”, they explicitly acknowledge these shortcuts. Such self-admitted technical debt (SATD) comments serve as signals [27, 20], marking where future maintenance is likely needed. Prior work has shown that SATD predicts maintenance challenges [18], guides refactoring decisions [37], and reveals developer concerns about code quality [36].
@@
As machine learning (ML) became widespread, researchers discovered new SATD patterns, including experimental pipelines and data dependencies, unlike traditional (non-ML) software, where technical debt primarily stems from structural and design issues [29, 3]. Since ChatGPT’s release in November 2022, Large Language Model (LLM) development has introduced fundamentally new architectural patterns that differ from both ML and traditional (non-ML) systems [26]. Modern LLM applications embed logic in prompts, use vector stores in Retrieval-Augmented Generation (RAG), coordinate multi-model workflows via agent frameworks, and depend on volatile external APIs [17, 7, 8]. These architectural shifts have introduced new forms of technical debt. For example, at the prompt layer, minor wording changes can break entire workflows [23], while orchestration frameworks may obscure critical business logic across distributed components [12]. At the infrastructure layer, token-based pricing can introduce economic debt that might compound with scale [28], and dependency on rapidly changing external APIs can create model lock-in risks [11]. Though practitioners encounter these issues daily [38], systematic empirical evidence remains limited.
@@
Most existing SATD research is based on data collected before the widespread adoption of LLMs, with datasets [27, 36, 3, 25] that do not capture modern LLM architectural patterns. For example, Bhatia et al. [3] conducted an empirical analysis of SATD in ML and non-ML systems and provided a taxonomy of such debts without considering LLM-based systems. In addition, even recent work on LLM technical debt [1] has examined only single-model prompt invocation and does not capture modern multi-agent, multi-model, or RAG-based workflows, nor does it examine the evolution of SATD over time or compare LLMs against ML and non-ML software systems. In this paper, we conduct the first empirical study of SATD in the LLM era. We replicate and extend Bhatia et al.’s study [3] by constructing a new curated dataset of 159 LLM repositories (developed since November 2022) and comparing them against 159 ML and 159 non-ML repositories from Bhatia et al.’s dataset, using the latest updates to their metadata, commit history, and code comments. We analyze over five million comment events to identify patterns in the introduction, persistence, and removal of SATD across various development paradigms.
@@
Our experiments show that, despite their architectural complexity, LLM repositories accumulate SATD at similar rates to ML ones (3.95% vs. 4.10%) but remain debt-free $2.4$ times longer than ML repositories (a median of 492 days vs. 204 days) before debt starts to accumulate rapidly. We identify three new SATD categories specific to LLMs, accounting for 22.4% of the classified debt: Model-Stack Workaround Debt (6.3%), Model Dependency Debt (11.2%), and Performance Optimization Debt (4.9%). When mapped to LLM development pipeline [13], we observe that debt is most concentrated in Deployment and Monitoring (30%) and Pretraining (23%) stages. Our analysis shows that foundational SATD (the first debt introduced in a file) is rarely removed (removal rates below 5%), with developers tending to clean more recent, localized debt while leaving foundational issues intact. The first SATD introduced into any file remains nearly permanent (only 4.2% ever removed in LLM repositories) despite 49.1% overall comment-level removal. Our model achieves 82% accuracy for LLM debt versus 96% for ML debt. This gap suggests that while ML debt remains highly predictable from traditional software metrics (confirming Bhatia et al. [3]), LLM debt involves new factors (prompt engineering, RAG orchestration, API volatility) that conventional commit-level features cannot fully capture.
This paper makes the following contributions:
@@
    We present the first empirical comparison of SATD in LLM, ML, and non-ML software, revealing that LLM development has unique debt accumulation patterns.
@@
    We extend the existing SATD taxonomy [3] by identifying three new technical debt types specific to LLM software.
@@
    We provide insights on where debt accumulates (deployment >> pretraining >> infrastructure), what predicts its persistence, and how debt removal differs across paradigms.
@@
    We release our complete replication package [30], including analysis scripts, our extended SATD detector, and the curated dataset of 159 LLM repositories with full git histories for replication and future studies.
@@ II Related Work
Technical Debt in Machine Learning Systems. ML systems introduce forms of technical debt that go beyond those found in traditional systems. Sculley et al. first highlighted how ML systems accumulate data dependency debt, configuration debt, and boundary erosion, noting that seemingly small design shortcuts can have widespread consequences through data pipelines and model behavior [29]. Building on this theoretical framework, Bhatia et al. [3] conducted a landmark empirical study comparing SATD in 318 ML and 318 non-ML repositories, finding that ML systems contain twice the SATD density of traditional (non-ML) software. They found ML repositories introduce debt $140$ times faster but also remove it $3.7$ times faster. Their work established the empirical foundation for understanding ML technical debt at scale. O’Brien et al. [25] further expanded this understanding by identifying 23 distinct types of ML-specific technical debt through qualitative analysis. Together, these studies underscore that ML systems carry structural, data-centric, and operational debt patterns that differ significantly from traditional systems, with Bhatia et al.’s quantitative findings providing the empirical basis that motivates our replication and extension to LLM systems.
Emerging Technical Debt in LLM-Based Development. Empirical research on LLM-specific SATD is still nascent. Aljohani and Do [1] examined SATD in early LLM API usage patterns and found that prompt design, rather than code structure, is a primary source of LLM-specific TD. However, their focus on single-model prompt invocation does not capture modern multi-agent, multi-model, or RAG-based workflows, nor does it examine the evolution of SATD over time.
@@ III Research Questions
This study replicates and extends Bhatia et al.’s [3] SATD analysis of ML systems to explore how SATD manifests in LLM systems. We examine whether architectural shifts in LLM development, such as prompt engineering, external APIs, and multi-model orchestration, change technical debt patterns compared to traditional ML and non-ML software. Accordingly, we compare the prevalence, types, stages, and survival of SATD in LLM, ML, and non-ML systems. Our study is structured around five research questions that mirror and extend those of Bhatia et al. [3].
@@
RQ2. What are the different types of SATD in LLM-based systems? The original study extended Bavota and Russo’s SATD taxonomy [2] to include ML-specific categories like configuration debt and inadequate testing. LLM development introduces new components absent in classical ML, such as RAG orchestration, vector databases, human-in-the-loop workflows, prompt templates, embedding models, token-cost optimization, context-window management, and agent orchestration. These elements may create new debt categories within complex multi-model pipelines. We explore whether LLM repositories exhibit distinct SATD patterns that require further taxonomy extensions beyond pre-LLM ML systems.
RQ3. Which stages of the LLM pipeline are more prone to SATD? Bhatia et al. [3] found that model building and data preprocessing stages accumulate the most debt in classical ML pipelines. We investigate whether this holds for LLM systems or if debt instead concentrates in new architectural components (prompt orchestration, agentic logic, and RAG/vector database integration) or in LLM-specific stages (like pretraining, fine-tuning, or deployment and monitoring).
@@ IV-A Replication Scope and Goals
To study SATD in LLM systems, we replicate and extend Bhatia et al.’s study of SATD in ML systems [3]. Their original analysis examined SATD in 318 ML and 318 non-ML Python projects using SATD detection and survival analysis to characterize SATD prevalence, types, and lifecycle. Our goals are twofold: Replication: Reproduce their main quantitative findings on SATD prevalence and survival in ML and non-ML repositories using their replication package, detector, and analysis pipeline, adapted to current software versions, as the original replication package required updates to deprecated library dependencies; Extension: Extend the study to contemporary, post-ChatGPT LLM repositories by constructing a curated dataset of LLM projects and applying the same SATD detection and survival analysis, as well as a refined SATD taxonomy tailored to LLM-specific development practices.
We follow the original study’s design wherever possible, including project curation criteria, comment extraction, SATD detection, and statistical tests, and clearly document deviations, such as the time span analyzed (temporal window), LLM-specific search queries, and taxonomy extensions.
@@ IV-B LLM Dataset Curation
Similar to Bhatia et al. [3], we collect balanced repository sets with complete project evolution history via GitHub API to retrieve active systems, following established MSR practices [14, 24]. We construct 41 search queries combining LLM-specific terms (“chatbot”, “gpt”, “openai”, “anthropic”), frameworks (“langchain”, “llama-index”), and architectural markers (“rag”, “embedding”, “agent”) with quality filters adapted from Bhatia et al.: (C1) non-forked repositories, (C2) $\geq 5\,$ Python files, (C3) >1>1 month of development history, and (C4) presence of LLM frameworks. Data collection was completed in early November 2025. A multi-signal classification pipeline confirmed LLM relevance through (1) README semantics, (2) metadata/topic labels, and (3) dependency inspection. We organize 159 LLM repositories into a two-layer taxonomy: Infrastructure/Tools (n=101) and End-User Applications (n=58), with functional tags including Evaluation/Testing, Agentic, Training/Fine-tuning, Serving/Inference, RAG, General Chatbot, and Prompt Engineering. Our dataset targets post-November 2022 repositories to capture contemporary LLM constructs (e.g., prompt engineering, RAG, agentic orchestration, API-driven invocation), distinct from Bhatia et al.’s pre-LLM dataset (2015-2021), enabling faithful SATD replication within the LLM era.
@@ IV-C ML and Non-ML Repository Selection
Though Bhatia et al. [3] analyzed 318 ML and 318 non-ML Python projects, we sample 159 repositories per category to match our LLM dataset as in Bavota and Russo [2]. Their work, which forms the basis for SATD classification in both Bhatia et al.’s and our study, showed that 159 repositories provide sufficient statistical power for SATD analysis while remaining computationally feasible for manual classification. This sample size ensures balanced comparison across LLM, ML, and non-ML repositories, without overrepresenting any type. For ML repositories, we randomly select 159 projects from Bhatia et al.’s replication package [^1] https://github.com/Aaditya-Bhatia/Self-Admitted-Technical-Debt-in-Machine-Learning-Software, ensuring coverage across their six domains (Image Processing, Natural Language Processing (NLP), Audio Processing, Autonomous Gameplay, Self-Driving Cars, Other ML). For non-ML repositories, we replicate their GitHub API methodology using the keywords ”python projects” and ”server” with identical filtering criteria (non-forked, $\geq$5 Python files, >>1 month history, no ML components), yielding 159 repositories that matched the ML dataset characteristics.
@@ IV-D SATD Detection Pipeline
We employ Liu et al.’s NLP-based SATD detector [19], a state-of-the-art SATD classification tool that achieved 0.82 F1-score in prior evaluations. The detection process involved three steps: (1) extracting Python comments with Comment-Parser, (2) applying the pre-trained classifier to identify SATD instances, and (3) tracking temporal metadata for survival analysis. For RQ1, we analyze SATD prevalence in the current codebase (502,592 unique comments in LLM repositories). For RQ4-RQ5, we track 3,713,429 comment events (additions and removals) across all commits in LLM repositories, enabling both cross-sectional prevalence assessment and longitudinal lifecycle tracking.
@@ IV-E SATD Sampling and Classification Procedure
To address RQ2, we select a representative subset of SATD comments from the dataset obtained in RQ1. Using stratified sampling across the two layers with a 95% confidence level and 5% margin of error, we obtain a random sample of 377 SATD instances. This ensures balanced domain representation and reduces bias toward any single application area. Before coding the sampled comments, two co-authors jointly examined an initial set of 70 SATD instances (distinct from the 377 used in the final analysis) to align on the classification scheme. Coding followed a consensus-based approach and a card-sorting process [33]: both authors discussed each instance, resolving uncertainties through collaborative negotiation. We report raw agreement: 60 of the 70 instances (85.7%) were coded without disagreement, while 10 instances (14.3%) required further discussion before reaching consensus. Through this iterative discussion, the two co-authors also identified patterns that did not fit existing categories and collectively defined three new LLM-specific debt types. This collaborative approach ensured consistent interpretation while allowing for emergent category discovery, which is particularly important given the novel characteristics of LLM-related technical debt.
After that, we categorize the 377 sampled SATD comments using a card-sorting approach guided by the extended taxonomy proposed by Bavota and Russo [2] [^2] Bhatia et al. [3] further adapted this taxonomy by adding Configuration Debt and Inadequate Testing, both of which we incorporate in our analysis.. Card sorting facilitated iterative comparison, refinement of category boundaries, and consistent labeling based on the initial joint examination. During coding, we encountered SATD comments that contained only the keyword “TODO” with little or no explanatory description of the intended action. We labeled such instances as Undefined, while those lacking context for technical debt were labeled as Unknown. Out of the 377 sampled instances, 44% (166) were identified as false positives. From the remaining 211 true SATD instances, we excluded those labeled as Undefined or Unknown , resulting in a final set of 143 relevant instances for RQ2. If an SATD instance did not fit the existing taxonomy, we introduce a new category, following Bhatia et al.’s approach of extending Hindle et al.’s taxonomy for ML systems [10].